Find the present value

What present value amount do you need to deposit in a bank account with 2.00% annual interest rate to have $100,000.00 at the end of 20 years?  
Answer #1

We use the formula:
A=P(1+r/100)^n
where
A=future value
P=present value
r=rate of interest
n=time period.

100,000=P*(1.02)^20

P=100,000/(1.02)^20

=$67297.13(Approx).